OpenStudy (anonymous):

What is the mass (in grams) of 9.66*10 to the 24 molecules on methanol (CH3OH)

OpenStudy (aaronq):

use these 2 equations: \(n=\dfrac{N}{N_A}\) ; n=moles, N=number of particles, \(N_A\) =avogadros number and \(n=\dfrac{m}{M}\) ; n=moles, M=molar mass (in g/mol), m=mass (in g)

OpenStudy (anonymous):

@aaronq but i dont know how i would set it up. i keep getting the wrong answer?

OpenStudy (anonymous):

can you plz help me solve it

OpenStudy (aaronq):

okay, in the first equation, you're looking for moles \(n=\dfrac{9.66*10^{24}}{6.022*10^{23}}\) then in the second equation you're looking for mass: \(n=\dfrac{9.66*10^{24}}{6.022*10^{23}}=\dfrac{m}{M_{methanol}}\) so look up the molar mass of methanol and solve for m

OpenStudy (anonymous):

im looking for it but im lost i cant find it:( @aaronq

OpenStudy (aaronq):

use the periodic table (you have to add the mass of the individual atoms), CH3OH is not an element

OpenStudy (anonymous):

okay so c its 12.01+1.01+16.00+1.01=30.03 so m = 30.03 is that correct

OpenStudy (anonymous):

@aaronq

OpenStudy (aaronq):

there are 4 hydrogens not 2

OpenStudy (aaronq):

12.01+(3)1.01+16.00+1.01

OpenStudy (anonymous):

@aaronq 32.05

OpenStudy (anonymous):

then wats the next step or thats the answer

OpenStudy (aaronq):

use the equation i wrote previously

OpenStudy (anonymous):

yeah but can you help me solve cuz i dont have a calculator and i dont know how plz

OpenStudy (anonymous):

@aaronq

OpenStudy (aaronq):

http://web2.0calc.com

OpenStudy (aaronq):

plug it into where molar mass goes and solve for m

OpenStudy (anonymous):

k i try

OpenStudy (anonymous):

@aaronq i cant:( it i cant do it can you solve it for me plz

OpenStudy (aaronq):

why can't you? show me what you did

OpenStudy (anonymous):

because in the example idk where 32.05 goes? so i divided by wat?

OpenStudy (aaronq):

M= molar mass \(M_{methanol}= 32.05\;g/mol\)

OpenStudy (anonymous):

so 32.05 g thats the answer? because i try it nad it was wrong?

OpenStudy (anonymous):

@aaronq

OpenStudy (aaronq):

dude, use a calculator. I'm willing to help you but not do it for you.

OpenStudy (anonymous):

aww u wont why?

OpenStudy (anonymous):

:(

OpenStudy (anonymous):

@abbieWhat is the mass (in grams) of 9.66*10 to the 24 molecules on methanol (CH3OH)

OpenStudy (anonymous):

can u help me plz @abbie

OpenStudy (abbie):

I'm not good at this but here's an example I found Divide 3.62x10^24 by 6.01 x 10^23 (Avogadro) Then you find about 6 mol methanol. The molar mass is (12 + 3 + 16 + 1) gram. you add them and get an answer.

OpenStudy (abbie):

work on it and tell me what you get

OpenStudy (anonymous):

@abbie yeah i did that and got 32.05 but it was it in grams

OpenStudy (anonymous):

im stuck

OpenStudy (abbie):

@amistre64 can you help him?

OpenStudy (anonymous):

@abbie what about this one how many atoms are in 187 g of calcium? _______atoms

OpenStudy (amistre64):

how much does one molecule of the stuff weigh?

OpenStudy (amistre64):

if m is equal to the mass of 1 molecule then: m * 9.66 * 10^(24), is the mass they are looking for right?

OpenStudy (anonymous):

YES

OpenStudy (anonymous):

i got that part i added the mass by individual and got 32.05 but then what??

OpenStudy (anonymous):

because i added C,H,O,H

OpenStudy (aaronq):

\(\dfrac{9.66∗10^{24}}{6.022∗10^{23}}=\dfrac{m}{32.05 g/mol}\rightarrow m=\dfrac{32.05 g/mol*9.66∗10^{24}molecules}{6.022∗10^{23}molecules/mol}\)

OpenStudy (anonymous):

so divided that or wat?] nad i got 5.32

OpenStudy (anonymous):

g

OpenStudy (anonymous):

@abbie how many atoms are in 187 g of calcium can u help do this one plz?

OpenStudy (aaronq):

you need to work on your algebra.

OpenStudy (anonymous):

i know and i got it wrong i havent had math is years thats why?

OpenStudy (anonymous):

can u help me solve it @aaronq

OpenStudy (aaronq):

you simply need to plug it into your calculator, it's 514.119893 g

OpenStudy (anonymous):

thanks . i have another question convert 0.108 mol of lithium to grams? ___g

OpenStudy (aaronq):

use \(n=\dfrac{m}{M}\)